# Phase error rate estimation in QKD with imperfect detectors

### v2

@@ VIII Summary and Discussion
In this work, we presented a finite-size security proof of QKD in the presence of imperfectly characterized and adversary controlled basis-efficiency mismatch. Thus, we addressed a longstanding assumption made in security proofs within the EUR and phase error correction frameworks. Before this work, proofs within these frameworks were not stable, and would be invalidated by infinitesimal amounts of basis-efficiency mismatch. Since our methods permit adversarial control over the efficiency mismatch, we also resolved a critical open problem of detector side-channels, which has remained unresolved in existing security proof approaches for standard QKD. Finally, we also fixed several technical issues in the security analysis of decoy-state QKD within the EUR framework.
@@
We applied our results to the decoy-state BB84 protocol, demonstrating practical key rates in the finite-size regime even in the presence of basis-efficiency mismatch. We also investigate quantitatively, the effect of methods such as random swapping of detectors, to reduce efficiency mismatch. Taken together, these results are a significant step towards protocol security of the EUR proof technique, and the implementation security of QKD using trusted detection setups.
@@
There are several opportunities for interesting work that builds upon these results. One possibility is to integrate our methods with established methods for addressing source imperfections [6, 7, 8]. Such a combination would lead to a security proof robust to both source and detector imperfections. Another avenue is applying these methods to passive detection setups, where the basis-efficiency mismatch assumption translates to an assumption of a perfectly balanced beam splitter and identical detectors. For computing key rates based on our results, suitable bounds on $\delta_{1},\delta_{2}$ are required. While our current analysis relies on some simplifications to obtain these bounds, they can likely be improved. Finally, an ambitious direction of future work would be to directly address correlated detectors. While our methods probably allow for one way of addressing this (Section VII.1), we find this unsatisfactory due to the requirement of protocol modifications.
@@ Appendix A Technical Statements
We use $S_{\circ}(Q)$ to denote the set of normalized states on $Q$. We use $S_{\bullet}(Q)$ to denote the set of all sub-normalized states on $Q$. We use $\text{Pos}(Q)$ to denote the set of positive semi-definite operators on $Q$. The smoothing on the min and max entropies is with respect to the purified distance [54, Definition 3.8].
@@
where we used the definition of the smooth max entropy in the first inequality, and [55, Sec. 4.3.2] for the second equality. The third inequality and the fourth equality follow from the definitions. The fifth inequality follows from the fact that the state $\tilde{\rho}$ is guaranteed to have $\leq ne_{\text{max}}$ errors, while the final inequality follows from the suitable bound on the sum of binomial coefficients. ∎
@@ B.0.1 Variable-length security definition
In order to prove the $\varepsilon_{\text{secure}}$-security of a variable-length QKD protocol [23], one must show that for all attacks by the adversary, the following statement is true :
@@ B.0.2 Bounding the smooth min entropy
We can make Bob measure his systems $B$ in the $X$ basis to obtain the classical outcome $X_{B}$. Then, using data processing [54, Theorem 6.2], we obtain
@@ B.0.3 Proving variable-length security
Here, we used the leftover-hashing lemma [2, Proposition 9] with the appropriate smoothing parameter on the sub-normalized state $(\rho_{|\Omega_{(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})}})_{\wedge\Omega_{\text{EV}}}$ for the first inequality. Since we use sub-normalized conditioning on $\Omega_{\text{EV}}$, it only appears in the smooth min entropy term and not inside the probability. Next, we use [2, Lemma 10] to get rid of the sub-normalized conditioning ($\wedge\Omega_{\text{EV}}$) in the smooth min entropy term in the second inequality. We used [54, Lemma 6.8] to split off the error-correction information ($\lambda_{\text{EC}}(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$) and error-verification information ($\log(2/\varepsilon_{\text{EV}})$) in the third inequality. We used the bound on the smooth min entropy from Eq. 78 for the fourth inequality, and the values of $l(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ and $\lambda_{\text{EC}}(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ from Eq. 6 for the fifth equality. We used concavity of the square root function and Jensen’s inequality to pull the sum over probabilities inside the square root for the sixth inequality, and Eq. 74 for the final inequality.
@@
Notice that in the variable-length protocol for which we proved security, the number of bits on which privacy amplification is applied is variable. It depends on the number of key generation rounds obtained in each protocol run. Some subtle issues regarding two-universal hashing on a variable-length input register were pointed out and addressed in [24, Section VII]. In particular, it was noted that first looking at the number of bits in the raw key, and then choosing an appropriate two-universal hashing procedure for that many input bits, does not produce a valid two-universal hashing procedure on the input space of variable-length bit strings. Due to this, the leftover-hashing lemma cannot be straightforwardly applied to such a scenario. However, this issue was addressed by showing that when the locations of the discard rounds are publicly announced, the theoretical analyses of scenarios where the rounds are actually discarded, vs mapped to special symbols such as 00 or $\bot$ (where leftover-hashing lemma can be applied), are equivalent [24, Lemmas 4, 5]. Due to this equivalence, the PA procedure described above can be applied in QKD protocols.
It is interesting to note that these issues are completely avoided by the above proof, in a very different manner than [24]. This is because in this proof, we always apply the leftover-hashing lemma on a state conditioned on the specific length of the raw key register (Eq. 78). Therefore, the leftover-hashing lemma can be applied in a straightforward manner, and there are no issues is choosing the hashing family based on the specific length of the raw key register. In other words, the PA procedure described above is valid for this proof. In a similar sense, the variable-length security proof from [24] critically relied on a technical lemma (Lemma 9), that necessitated the use of Rényi entropies instead of smooth min entropy in that work. However, the variable-length proof presented above takes a different approach, and does not impose the same requirements on the behaviour of smooth min entropy. Again, this is due to the use of Eq. 78.
@@ C.2 Sampling with imperfect detectors
$\Pr(\frac{\bm{N}_{P^{\prime}}}{n}\geq e+(\delta+c))\leq\Pr(\frac{\bm{N}_{P}}{n}\geq e)+F(n,c,\delta),$ (86)
where $F(n,c,\delta)$ was defined in Lemma 4.
@@
The proof follows after noting that Eqs. 87 and 88 $\implies$ Eq. 86. ∎
@@
$\Pr(\bm{N}_{G^{\prime}+\delta I}\geq ne+n(2\delta+c))\leq\Pr(\bm{N}_{G}\geq ne)+F(n,c,2\delta),$ (90)
@@
The proof follows from the observation that Eqs. 89, 90 and 91 $\implies$ Eq. 24. ∎
@@ C.3 Sampling with independent imperfect detectors
$\Pr(\frac{\bm{N}_{P^{\prime}}}{n}\geq e+(\delta+c))\leq\Pr(\frac{\bm{N}_{P}}{n}\geq e)+F(n,c,\delta),$ (92)
where $F(n,c,\delta)$ was defined in Lemma 4.
@@
$\Pr(\frac{\bm{N}_{P^{\prime}}}{n}\geq e+2\delta+c)\leq\Pr(\frac{\bm{N}_{P}}{n}\geq e)+F(n,c,2\delta),$ (93)
where $F(n,c,\delta)$ was defined in Lemma 4.
@@ Appendix E Decoy Analysis
Consider all the rounds where $O$ is observed. Condition on the event that $n_{O}$ such rounds are observed. Let $X_{1}\dots X_{n_{O}}$ be the sequence of photon number of Alice’s signal corresponding to these round. Condition further on the event that a specific sequence $X_{1}\dots X_{n_{O}}$ is observed.
@@ F.0.1 Bounding the smooth min entropy
$\Pr(\bm{e^{\text{key}}_{X,{1}}}\geq\mathcal{B}_{e}(\bm{e^{\text{obs}}_{X,\mu_{\vec{k}}}},\bm{n_{X,\mu_{\vec{k}}}},\bm{n_{K,\mu_{\vec{k}}}})\quad\lor\quad\bm{n_{K,{1}}}\leq\mathcal{B}_{1}(\bm{n_{K,\mu_{\vec{k}}}}))_{\Omega_{(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z},n_{K,{1}})}},$
@@
$\geq n_{K,{1}}(1-\mathcal{B}_{e}(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z}))$
where we used [54, Lemma 6.7] to isolate the single-photon contribution to the key in the first inequality, and the entropic uncertainty relations [1] followed by the same series of steps as in the proof of Theorem 1 to replace the smooth max entropy term with the bound on the phase error rate. This is the required bound on the smooth min entropy of the raw key.
@@
If one wishes to obtain key from multi-photon or zero-photon events, one can use a chain rules on the smooth min entropy at this stage (as is done in [30]). However, note that one must first fix the number of pulses corresponding to these events in order to have the registers be well-defined. For instance, one cannot apply the above analysis for a state $\rho$ without conditioning on a $\widetilde{\Omega}$, since a fixed value of $n_{K,{1}}$ is required to meaningfully define the register $Z^{n_{K,{1}}}$. This subtlety is missing in [30]. While one may choose to define these registers to store strings of variable length, this then complicates the application of EUR in the proof (since the number of rounds on which the EUR is applied is now variable). Our analysis is one rigorous way to avoid these problems.
@@
$\Delta$ $=\frac{1}{2}\sum_{\begin{subarray}{c}n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z}\\ n_{K,{1}}\geq\mathcal{B}_{1}(n_{K,\mu_{\vec{k}}})\end{subarray}}\Pr(\widetilde{\Omega}\wedge\Omega_{\text{EV}})d(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z},n_{K,{1}})$ (107)
@@
$\Delta_{1}$ $\leq\sum_{\begin{subarray}{c}n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z}\in\mathcal{F}\\ n_{K,{1}}\geq\mathcal{B}_{1}(n_{K,\mu_{\vec{k}}})\end{subarray}}\Pr(\widetilde{\Omega})\bigg{(}2\sqrt{\kappa(\widetilde{\Omega})}+\frac{1}{2}2^{-\frac{1}{2}\left(H^{\sqrt{\kappa(\widetilde{\Omega})}}_{\text{min}}(Z^{n_{K}}|E^{n}C^{n}C_{E})_{(\rho|\widetilde{\Omega})\wedge\Omega_{\text{EV}}}-l(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z})\right)}\bigg{)}$ (109)
$\leq\sum_{\begin{subarray}{c}n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z}\in\mathcal{F}\\ n_{K,{1}}\geq\mathcal{B}_{1}(n_{K,\mu_{\vec{k}}})\end{subarray}}\Pr(\widetilde{\Omega})\bigg{(}2\sqrt{\kappa(\widetilde{\Omega})}+\frac{1}{2}2^{-\frac{1}{2}\left(H^{\sqrt{\kappa(\widetilde{\Omega})}}_{\text{min}}(Z^{n_{K}}|E^{n}C^{n}C_{E})_{\rho|\widetilde{\Omega}}-l(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z})\right)}\bigg{)}$
$\leq\sum_{\begin{subarray}{c}n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z}\in\mathcal{F}\\ n_{K,{1}}\geq\mathcal{B}_{1}(n_{K,\mu_{\vec{k}}})\end{subarray}}\Pr(\widetilde{\Omega})\bigg{(}2\sqrt{\kappa(\widetilde{\Omega})}$

### v3

@@ IX Summary and Discussion
In this work, we presented a finite-size security proof of the decoy-state BB84 protocol in the presence of imperfectly characterized and (bounded) adversary controlled basis-efficiency mismatch. Thus, we addressed a longstanding assumption made in security proofs for such protocols within the EUR and phase error correction frameworks. Before this work, proofs within these frameworks were not stable, and would be invalidated by infinitesimal amounts of basis-efficiency mismatch (This problem does not arise in MDI-QKD, and is not resolved by this work for entanglement-based protocols). Since our methods permit (bounded) adversarial control over the efficiency mismatch, we also develop a framework to address an important class of detector side-channels, which has remained unresolved in existing security proof approaches for standard QKD.
@@
We also fixed several technical issues in the security analysis of decoy-state QKD within the EUR framework. We applied our results to the decoy-state BB84 protocol, demonstrating practical key rates in the finite-size regime even in the presence of basis-efficiency mismatch. We also investigate quantitatively, the effect of methods such as random swapping of detectors, to reduce efficiency mismatch. Taken together, these results are a significant step towards protocol security of the EUR proof technique, and the implementation security of QKD using trusted detection setups.
Additionally, we sketch out a possible approach towards handling correlated detectors — an unresolved problem with little progress so far. Moreover, although the rigorous results we obtain for multi-mode detectors depend on a specific model, we expect this framework to be adapted to more realistic models with subsequent work.
@@
There are several opportunities for interesting work that builds upon these results. One natural extension is to integrate our methods with established methods for addressing source imperfections [6, 7, 8]. Such a combination would lead to a security proof robust to both source and detector imperfections. Another avenue is applying these methods to passive detection setups, where the basis-efficiency mismatch assumption translates to an assumption of a perfectly balanced beam splitter and identical detectors. For computing key rates based on our results, suitable bounds on $\delta_{1},\delta_{2}$ are required. While our current analysis relies on some simplifications to obtain these bounds, they can likely be improved.
Moreover, while we present a possible approach in this work, handling correlated detectors in a fully rigorous manner is still an open task. Finally, examining a wider spectrum of detector imperfections – well characterized by state-of-the-art experimental methods – would further broaden the applicability of our results. Such an endeavor would require close collaboration with experimentalists to refine the characterization of imperfections, as well as theoretical advancements to extend the framework to encompass a wider class of side channels.
Acknowledgements
We thank Victor Zapatero, Margarida Pereira, and Guillermo Currás-Lorenzo for useful discussions on the phase error correction framework and the EUR approach to QKD security proofs, and for providing helpful comments on this manuscript. We thank Ernest Tan for useful discussions that led to the variable-length security proof in this work. We thank Ashutosh Marwah for some helpful comments on our random sampling arguments. We thank Masato Koashi, and Antia Lamas-Linares for helpful discussions on the detector model used in this work. We thank John Burniston for help with code. We thank Federico Grasselli for detailed feedback on early versions of this work. This work was funded by the NSERC Discovery Grant, and was conducted at the Institute for Quantum Computing, University of Waterloo, which is funded by the Government of Canada through ISED. DT and PS are partially funded by the Mike and Ophelia Laziridis Fellowship.
@@ Appendix A Technical Statements
We use $S_{\circ}(Q)$ to denote the set of normalized states on $Q$. We use $S_{\bullet}(Q)$ to denote the set of all sub-normalized states on $Q$. We use $\text{Pos}(Q)$ to denote the set of positive semi-definite operators on $Q$. The smoothing on the min and max entropies is with respect to the purified distance [65, Definition 3.8].
@@
where we used the definition of the smooth max entropy in the first inequality, and [66, Sec. 4.3.2] for the second equality. The third inequality and the fourth equality follow from the definitions. The fifth inequality follows from the fact that the state $\tilde{\rho}$ is guaranteed to have $\leq ne_{\text{max}}$ errors, while the final inequality follows from the suitable bound on the sum of binomial coefficients. ∎
@@ B.0.1 Variable-length security definition
In order to prove the $\varepsilon_{\text{secure}}$-security of a variable-length QKD protocol [25, 26], one must show that for all attacks by the adversary, the following statement is true :
@@ B.0.2 Bounding the smooth min entropy
We can make Bob measure his systems $B$ in the $X$ basis to obtain the classical outcome $X_{B}$. Then, using data processing [65, Theorem 6.2], we obtain
@@ B.0.3 Proving variable-length security
Here, we used the leftover-hashing lemma [2, Proposition 9] with the appropriate smoothing parameter on the sub-normalized state $(\rho_{|\Omega_{(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})}})_{\wedge\Omega_{\text{EV}}}$ for the first inequality. Since we use sub-normalized conditioning on $\Omega_{\text{EV}}$, it only appears in the smooth min entropy term and not inside the probability. Next, we use [2, Lemma 10] to get rid of the sub-normalized conditioning ($\wedge\Omega_{\text{EV}}$) in the smooth min entropy term in the second inequality. We used [65, Lemma 6.8] to split off the error-correction information ($\lambda_{\text{EC}}(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$) and error-verification information ($\log(2/\varepsilon_{\text{EV}})$) in the third inequality. We used the bound on the smooth min entropy from Eq. 79 for the fourth inequality, and the values of $l(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ and $\lambda_{\text{EC}}(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ from Eq. 6 for the fifth equality. We used concavity of the square root function and Jensen’s inequality to pull the sum over probabilities inside the square root for the sixth inequality, and Eq. 75 for the final inequality.
@@
Notice that in the variable-length protocol for which we proved security, the number of bits on which privacy amplification is applied is variable. It depends on the number of key generation rounds obtained in each protocol run. Some subtle issues regarding two-universal hashing on a variable-length input register were pointed out and addressed in [27, Section VII]. In particular, it was noted that first looking at the number of bits in the raw key, and then choosing an appropriate two-universal hashing procedure for that many input bits, does not produce a valid two-universal hashing procedure on the input space of variable-length bit strings. Due to this, the leftover-hashing lemma cannot be straightforwardly applied to such a scenario. However, this issue was addressed by showing that when the locations of the discard rounds are publicly announced, the theoretical analyses of scenarios where the rounds are actually discarded, vs mapped to special symbols such as 00 or $\bot$ (where leftover-hashing lemma can be applied), are equivalent [27, Lemmas 4, 5]. Due to this equivalence, the PA procedure described above can be applied in QKD protocols.
It is interesting to note that these issues are completely avoided by the above proof, in a very different manner than [27]. This is because in this proof, we always apply the leftover-hashing lemma on a state conditioned on the specific length of the raw key register (Eq. 79). Therefore, the leftover-hashing lemma can be applied in a straightforward manner, and there are no issues is choosing the hashing family based on the specific length of the raw key register. In other words, the PA procedure described above is valid for this proof. In a similar sense, the variable-length security proof from [27] critically relied on a technical lemma (Lemma 9), that necessitated the use of Rényi entropies instead of smooth min entropy in that work. However, the variable-length proof presented above takes a different approach, and does not impose the same requirements on the behavior of smooth min entropy. Again, this is due to the use of Eq. 79.
@@ C.2 Sampling with imperfect detectors
$\Pr(\frac{\bm{N}_{P^{\prime}}}{n}\geq e+(\delta+c))\leq\Pr(\frac{\bm{N}_{P}}{n}\geq e)+F(n,\delta,c),$ (87)
where $F(n,\delta,c)$ was defined in Lemma 4.
@@
The proof follows after noting that Eqs. 88 and 89 $\implies$ Eq. 87. ∎
@@
$\Pr(\bm{N}_{G^{\prime}+\delta I}\geq ne+n(2\delta+c))\leq\Pr(\bm{N}_{G}\geq ne)+F(n,2\delta,c),$ (91)
@@
The proof follows from the observation that Eqs. 90, 91 and 92 $\implies$ Eq. 24. ∎
@@ C.3 Sampling with independent imperfect detectors
$\Pr(\frac{\bm{N}_{P^{\prime}}}{n}\geq e+(\delta+c))\leq\Pr(\frac{\bm{N}_{P}}{n}\geq e)+F(n,\delta,c),$ (93)
where $F(n,\delta,c)$ was defined in Lemma 4.
@@
where $F(n,\delta,c)$ was defined in Lemma 4.
@@ Appendix E Decoy Analysis
Consider all the rounds where $O$ is observed. Condition on the event that $n_{O}$ such rounds are observed. Let $X_{1}\dots X_{n_{O}}$ be the sequence of photon numbers of Alice’s signals corresponding to these rounds. Condition further on the event that a specific sequence $X_{1}\dots X_{n_{O}}$ is observed.
@@ F.0.1 Bounding the smooth min entropy
$\Pr(\bm{e^{\text{key}}_{X,{1}}}\geq\mathcal{B}_{e}(e^{\text{obs}}_{X,\mu_{\vec{k}}},n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}})\quad\lor\quad\bm{n_{K,{1}}}\leq\mathcal{B}_{1}(n_{K,\mu_{\vec{k}}}))_{\Omega_{(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z},n_{K,{1}})}},$
@@
$\geq n_{K,{1}}\left(1-\mathcal{B}_{e}\left(h(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z})\right)\right)$
where we used [65, Lemma 6.7] to isolate the single-photon contribution to the key in the first inequality, and the entropic uncertainty relations [1] followed by the same series of steps as in the proof of Theorem 1 to replace the smooth max entropy term with the bound on the phase error rate. This is the required bound on the smooth min entropy of the raw key.
@@
If one wishes to obtain key from multi-photon or zero-photon events, one can use a suitable chain rule on the smooth min entropy at this stage (as is done in [35]). However, note that one must first fix the number of pulses corresponding to these events in order to have the registers be well-defined. For instance, one cannot apply the above analysis for a state $\rho$ without conditioning on a $\widetilde{\Omega}$, since a fixed value of $n_{K,{1}}$ is required to meaningfully define the register $Z^{n_{K,{1}}}$. This subtlety is missing in [35]. While one may choose to define these registers to store strings of variable length, this then complicates the application of EUR in the proof (since the number of rounds on which the EUR is applied is now variable). Our analysis is one rigorous way to avoid these problems.
@@
$\Delta$ $=\frac{1}{2}\sum_{\begin{subarray}{c}n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z}\\ n_{K,{1}}>\mathcal{B}_{1}(n_{K,\mu_{\vec{k}}})\end{subarray}}\Pr(\widetilde{\Omega}\wedge\Omega_{\text{EV}})d(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z},n_{K,{1}})$ (108)
@@
$\Delta_{1}$ $\leq\sum_{\begin{subarray}{c}n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z}\in\mathcal{F}\\ n_{K,{1}}>\mathcal{B}_{1}(n_{K,\mu_{\vec{k}}})\end{subarray}}\Pr(\widetilde{\Omega})\bigg{(}2\sqrt{\kappa(\widetilde{\Omega})}+\frac{1}{2}2^{-\frac{1}{2}\left(H^{\sqrt{\kappa(\widetilde{\Omega})}}_{\text{min}}(Z^{n_{K}}|E^{n}C^{n}C_{E})_{(\rho|\widetilde{\Omega})\wedge\Omega_{\text{EV}}}-l(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z})\right)}\bigg{)}$ (110)
$\leq\sum_{\begin{subarray}{c}n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z}\in\mathcal{F}\\ n_{K,{1}}>\mathcal{B}_{1}(n_{K,\mu_{\vec{k}}})\end{subarray}}\Pr(\widetilde{\Omega})\bigg{(}2\sqrt{\kappa(\widetilde{\Omega})}+\frac{1}{2}2^{-\frac{1}{2}\left(H^{\sqrt{\kappa(\widetilde{\Omega})}}_{\text{min}}(Z^{n_{K}}|E^{n}C^{n}C_{E})_{\rho|\widetilde{\Omega}}-l(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z})\right)}\bigg{)}$
$\leq\sum_{\begin{subarray}{c}n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z}\in\mathcal{F}\\ n_{K,{1}}>\mathcal{B}_{1}(n_{K,\mu_{\vec{k}}})\end{subarray}}\Pr(\widetilde{\Omega})\bigg{(}2\sqrt{\kappa(\widetilde{\Omega})}$